OpenStudy (anonymous):

Don't get this step. x^2+y^2=3y. then it says x^2+(y-3/2)^2= (3/2)^2 . where does the 3/2 come from?

OpenStudy (anonymous):

I should mention this is finding a Cartesian equation. r=3sin(theta)

OpenStudy (anonymous):

Took the 3y to the other side and complate that square.

myininaya (myininaya):

they seen it was a cirlce so they wrote it in the standard from of a circle

myininaya (myininaya):

x^2+y^2-3y=0 (x-0)^2+y^2-3y+(3/2)^2=(3/2)^2 (x-0)^2+(y-3/2)^2=(3/2)^2

OpenStudy (anonymous):

...This Calculus thing would be a whole lot easier if I ever actually learned Algebra. Thanks guys!